OpenStudy (anonymous):

Help please!!!! Rotated Ellipse with the equation x^(2)+xy+y^(2)=1. Find all the points where the tangent line is horizontal. Then find all points where the tangent line is parallel to the line y=-x.

OpenStudy (freckles):

so first question find for what x values we have y'=0 second question find for what x values we have the y'=-1

OpenStudy (freckles):

or all (x,y) such that y'=0 (question 1) and find all (x,y) such that y'=-1 (question 2)

OpenStudy (freckles):

have you found y' yet?

OpenStudy (anonymous):

Yeah, y'= (-2x-y)/(x+2y)

OpenStudy (freckles):

\[\text{ so } y'=0 \text{ when } -2x-y=0 \\ \text{ and } y'=-1 \text{ when } \frac{-2x-y}{x+2y}=-1\]

OpenStudy (freckles):

It is easy to solve that first equation for y it is a little more difficult to solve that bottom equation for y but not impossible

OpenStudy (freckles):

also you should probably exclude when x+2y can be zero from both answers

OpenStudy (anonymous):

when you get y=-2x what do you do to get the points that make it horizontal? And I'm not sure how to go about solving the second one. :/

OpenStudy (anonymous):

hey, you found y = -2x but we want the points on the ellipse so we plug y=-2x to the ellipse equation: x^2 + x * (-2x) + (-2x)^2 = 1 x^2 -2x^2 +4x^2 = 1 x^2 = 1/3 so you have two values of x..

OpenStudy (anonymous):

So +/- sqrt(1/3)?

OpenStudy (anonymous):

yes x values of the points so you can find now the y values as well

OpenStudy (anonymous):

what would you do for that?

OpenStudy (anonymous):

you can use either the ellipse equation or y=-2x

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

do you understand why we have to plug y=-2x into the ellipse equation ?

OpenStudy (anonymous):

To find the points that make the tangent line horizontal.

OpenStudy (anonymous):

when we took the derivative of the ellipse we got y=-2x. now since this derivative may be the same for different ellipses (for example with other values then 1 in the right side) we couldnt find the points coordinates. we had to put this result in the specific ellipse equation

OpenStudy (anonymous):

i hope it make sense

OpenStudy (anonymous):

To find the points where the tangent line is parallel to the line y=-x what would you do??

OpenStudy (anonymous):

use the derivative and find a relation between y and x then again plug it into the ellipse equation

OpenStudy (anonymous):

the slope is -1 would you set the derivative equal to that?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

is that x=-y?

OpenStudy (anonymous):

i think y=x

OpenStudy (anonymous):

(-2x-y)/(x+2y) = -1 -2x-y=-x-2y 2y-y=2x-x y=x

OpenStudy (anonymous):

then just plug that into the ellipse equation?

OpenStudy (anonymous):

thanks for your help! :)

OpenStudy (anonymous):

yes then you will find values of x\y and then you will find the points

OpenStudy (anonymous):

sure no problem

OpenStudy (anonymous):

is it x=1/3 again?

OpenStudy (anonymous):

well, the sqrt1/3

OpenStudy (anonymous):

so then the points would have the same x and y values?

OpenStudy (anonymous):

yes i think so, but you will get different y values

OpenStudy (anonymous):

wouldn't they be the same because y=x?

OpenStudy (anonymous):

sorry i meant different from the first case y'=0

OpenStudy (anonymous):

for y'=-1 the x and y are the same indeed

OpenStudy (anonymous):

(sqrt1/3,sqrt1/3)(-sqrt1/3,-sqrt1/3)?

OpenStudy (anonymous):

yes :) the points for y'=-1

OpenStudy (anonymous):

Okay! Thanks again.

OpenStudy (anonymous):

yw